OpenStudy (anonymous):

C 8 4 I need help solving this. please

OpenStudy (anonymous):



OpenStudy (anonymous):

I know that n is 8 and r is 4 I also know that 8-4=4

OpenStudy (anonymous):

and I know that the top of the equation is 8*7*6*5*4

OpenStudy (anonymous):

how do I figure out the rest?

OpenStudy (anonymous):

\[^{n}C _{r} = n!\space/(n-r)!\space*\space r!\] \[^{8}C _{4} = 8!\space/(8-4)!\space*\space 4!\] \[^{8}C _{4} = 8!\space/4!\space*\space 4!\] \[^{8}C _{4} = 8*7*6*5*4!\space/4!\space*\space 4!\] \[^{8}C _{4} = 8*7*6*5\space/ 4!\]

OpenStudy (anonymous):

I am so lost! I have been trying this for over an hour and I am getting no were

OpenStudy (anonymous):

wow your good! I would have never got that! Thank you

OpenStudy (anonymous):

Anytime. :)

OpenStudy (anonymous):

so the answer is 420 correct?

OpenStudy (anonymous):

No the answer is 70! pellet I am still lost!

OpenStudy (anonymous):

70 is the correct

OpenStudy (anonymous):

how? I keep getting 420 or 160 I suck

OpenStudy (anonymous):

8*7*6*5 = 1680 4*3*2*1=24 1680/24 = 70

OpenStudy (anonymous):

wow ok Thank you again. Now I feel even more dumb then before.

OpenStudy (anonymous):

You are just confused between topics. Nothing else.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

anytime :)